# bibliography.bib

### v1

@@ 4 Extensions of frame maps
Suppose that we have a frame map $f:L\rightarrow M$ and Raney extensions $(L,C)$ and $(M,D)$. If the map $\mathsf{cl}_{D^{*}}(f[-]):C^{*}\to D^{*}$ preserves all meets, then it is a coframe map.
@@
It suffices to observe that we are in the situation described by Lemma 4.1. This is because the map $\mbox{cl}_{D^{*}}(f[-])$ lifts a frame map $f:L\rightarrow M$, and so its restriction to $L$ is the composition of the join-preserving map $f$ with the join-preserving inclusion $M\subseteq D^{*}$. ∎
For a poset map to preserve all meets, it suffices for it to have a left adjoint. In general, for all $F\in C^{*}$ and $G\in D^{*}$, we have $\mbox{cl}_{D^{*}}(f[F])\subseteq G$ if and only if $f[F]\subseteq G$, and this holds if and only if $F\subseteq f^{-1}(G)$. As the order on $C^{*}$ and $D^{*}$ is reverse set inclusion, if a well-defined preimage map $f^{-1}:D^{*}\to C^{*}$ exists, then it is the left adjoint we are looking for. A frame map $f:L\to M$ always determines a preimage (frame) map $f^{-1}:\mathsf{Filt}(M)\to\mathsf{Filt}(L)$. The issue with the preimage is that it is not guaranteed to map filters of $D^{*}$ to filters of $C^{*}$.
@@
Suppose that $f:L\rightarrow M$ is a frame map and that we have Raney extensions $(L,C)$ and $(M,D)$. The map $f$ can be extended to a map of these Raney extensions if and only if $f^{-1}(G)\in C^{*}$ for every $G\in D^{*}$.
@@
It remains to show that, if a left adjoint of $\mbox{cl}_{D^{*}}(f[-])$ exists, then it has to be the preimage map. For a left adjoint $\mathcal{L}$ and for each $F\in C^{*}$ and $G\in D^{*}$, by definition of the closure $\mbox{cl}_{D^{*}}$, we would have $F\subseteq\mathcal{L}(G)$ if and only if $f[F]\subseteq G$. For each $x\in L$, we have $x\in f^{-1}(G)$ if and only if $f(x)\in G$, and this is in turn equivalent to $f[{\uparrow}x]\subseteq G$, as noticed above this holds if and only if ${\uparrow}x\subseteq\mathcal{L}(G)$, that is $x\in\mathcal{L}(G)$. For the last step, we have used the fact that ${\uparrow}x\in C^{*}$, as this collection is assumed to contain all principal filters of $L$. We have then shown that, for each $x\in L$, and each $G\in D^{*}$, $x\in f^{-1}(G)$ if and only if $x\in\mathcal{L}(G)$. ∎
We also notice that for a frame map $f:L\to M$ which can be extended to a Raney map $(L,C)\to(M,D)$, the definition of the extension as a coframe map $C\to D$ has to be $c\mapsto\bigwedge\{f(a):a\in{\uparrow}^{L}c\}$, from the fact that it has to extend $f$, together with preservation of all meets. We apply Theorem 4.3 above to some of the Raney extensions that we have seen.
@@
Any frame morphism $f:L\rightarrow M$ lifts to a coframe morphism
$f_{\mathcal{SE}}:\mathsf{Filt}_{\mathcal{SE}}(L)^{op}\rightarrow\mathsf{Filt}_{\mathcal{SE}}(M)^{op}.$
@@
By Theorem 4.3, a frame morphism $f:L\rightarrow M$ lifts as required if preimages of strongly exact filters are strongly exact. Suppose, then, that $F\subseteq M$ is strongly exact. Suppose that the meet $\bigwedge_{i}x_{i}$ is strongly exact, and that $f(x_{i})\in F$. Because all frame morphisms preserve strongly exact meets, as well as strong exactness of meets, we have $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})\in F$, as desired. ∎
Corollary 4.5.
Every map of frames $L\to M$ lifts to a coframe map $\mathsf{S}_{\mathfrak{o}}(L)\to\mathsf{S}_{\mathfrak{o}}(M)$.

### v2

@@ 5.2 The TD axiom
A $T_{0}$ space is $T_{D}$ if and only if all neighborhood filters are exact.
@@
Suppose that $X$ is a $T_{D}$ topological space. Neighborhood filters are completely prime, and by Proposition 2.17 all primes of the form $X{\setminus}\overline{\{x\}}$ are covered. Hence, by the characterization in Proposition 4.12, the corresponding neighborhood filters are exact. Conversely, if $X$ is not $T_{D}$ there must be a point $x\in X$ whose prime is not covered, and by Proposition 4.12 again, this means that its completely prime filter is not exact. ∎
@@
The following are equivalent for a $T_{0}$ space $X$.
  1. 1.
    The space $X$ is $T_{D}$.
  2. 2.
    The Raney extension $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{E}$-dense.
  3. 3.
    The Raney extension $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{E}$-canonical.
  4. 4.
    The Raney extension $(\Omega(X),\mathcal{U}(X))$ is isomorphic to $(\Omega(X),\mathsf{Filt}_{\mathcal{E}}(\Omega(X)))$.
  5. 5.
    The inclusion $\Omega(X)\subseteq\mathcal{U}(X)$ preserves exact meets.
@@
Let $X$ be a $T_{0}$ space. If this is a $T_{D}$ space, then by Lemma 5.6 all neighborhood filters are exact, and this means that all filters of the form ${\uparrow}^{\Omega(X)}{\uparrow}x$ for $x\in X$ are exact. As, for all $x\in X$, $\bigcap{\uparrow}^{\Omega(X)}{\uparrow}x={\uparrow}x$, and the principal filters generate the collection $\mathcal{U}(X)$, (2) follows. Suppose, now, that (2) holds. By Proposition 3.8, and the characterization of compactness of Proposition 3.2, the Raney extension is $\mathcal{E}$-compact, hence $\mathcal{E}$-canonical by our initial hypothesis. Items (3) and (4) are equivalent by the uniqueness result of Theorem 3.6. Suppose that (4) holds. We will identify $(\Omega(X),\mathcal{U}(X))$ with the isomorphic Raney extension $(\Omega(X),\mathsf{Filt}_{\mathcal{E}}(\Omega(X))^{op})$. If $U_{i}\in\Omega(X)$ is a family such that their meet is exact, then the least upper bound of the family ${\uparrow}U_{i}$ in $\mathsf{Filt}_{\mathcal{E}}(\Omega(X))$ must be ${\uparrow}\bigwedge_{i}U_{i}$, by definition of exact filter. This means that the meet is preserved by the embedding $\Omega(X)\to\mathsf{Filt}_{\mathcal{E}}(\Omega(X))^{op}$. Finally, (5) implies (1) by the characterization in Theorem 2.10. ∎
Motivated by the last result, we call a Raney extension $T_{D}$ if it is $\mathcal{E}$-dense. All Raney extensions are $\mathcal{E}$-compact, by Proposition 3.8 and the characterization of compactness in Proposition 3.2. Thus, the $T_{D}$ Raney extensions are those which are $\mathcal{E}$-canonical, and by the uniqueness result of Theorem 3.6 these are the Raney extensions which are, up to isomorphism, $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ for some frame $L$. In Section 6, we will study $T_{D}$ Raney extensions.
5.3 The T1 axiom
Let us now look at the $T_{1}$ axiom. The axiom $T_{1}$, too, can be characterized in terms of filters.
@@ 5.3 The T1 axiom
A $T_{0}$ space is $T_{1}$ if and only if all its neighborhood filters are regular.
@@
Suppose that $X$ is a $T_{1}$ space, and let $x\in X$. As $X$ is $T_{1}$, the set $X{\setminus}\{x\}$ is open. Then $N(x)=\{U\in\Omega(X):U\cup(X{\setminus}\{x\})=X\}$. By the characterization of regular filters in Proposition 2.12, this is a regular filter. For the converse, suppose that $X$ is a $T_{0}$ space where all neighborhood filters are regular. Let $x\in X$. We will show that $\{x\}$ is closed by showing ${\downarrow}x=\{x\}$. By the characterization in Proposition 2.12, and because neighborhood filters are completely prime, there is some open $V\in\Omega(X)$ such that:
$N(x)=\{U\in\Omega(X):U\cup V=X\}.$
Observe that $\bigcap N(x)\cup V={\uparrow}x\cup V=X$, thus $V^{c}\subseteq{\uparrow}x$. Since $\emptyset\notin N(x)$, $V\cup P=\emptyset\cup V\neq X$. Then, $V\notin N(x)$ and so $x\in V^{c}$. As $V^{c}$ is a downset in the specialization order, ${\downarrow}x\subseteq V^{c}$. But this means ${\downarrow}x\subseteq V^{c}\subseteq{\uparrow}x$, hence ${\downarrow}x=\{x\}$. ∎
Lemma 5.9.
If $(\Omega(X),\mathcal{P}(X))$ is a Raney extension then it is $(\Omega(X),\mathsf{Filt}_{\mathcal{R}}(\Omega(X))^{op})$, up to isomorphism.
Proof.
Let $X$ be a space. We observe that if $(\Omega(X),\mathcal{P}(X))$ is a Raney extension, each subset is an intersection of opens, by meet generativity, and so $X$ is $T_{1}$. To show our claim, it suffices to show that $\mathcal{P}(X)^{*}=\mathsf{Filt}_{\mathcal{R}}(\Omega(X))$. Filters in $\mathcal{P}(X)^{*}$ are those of the form ${\uparrow}^{\Omega(X)}S$ for arbitrary subsets $S\subseteq X$. As $X$ is $T_{1}$, $S$ is closed, so for each $S\subseteq X$, ${\uparrow}^{\Omega(X)}S=\{U\in\Omega(X):S^{c}\cup U=X\}$. Since $S^{c}$ is open, this filter is regular by Proposition 2.12. ∎
Theorem 5.10.
The following are equivalent for a $T_{0}$ space $X$.
  1. 1.
    The space $X$ is $T_{1}$.
  2. 2.
    The pair $(\Omega(X),\mathcal{P}(X))$ is a Raney extension, and equals $(\Omega(X),\mathcal{U}(X))$.
  3. 3.
    The Raney extension $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{R}$-dense.
  4. 4.
    The Raney extension $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{R}$-canonical.
  5. 5.
    The Raney extension $(\Omega(X),\mathcal{U}(X))$ is isomorphic to $(\Omega(X),\mathsf{Filt}_{\mathcal{R}}(\Omega(X)))$.
Proof.
The equivalence between (1) and (2) holds as $\Omega(X)$ meet-generates $\mathcal{P}(X)$ if and only if $X$ is $T_{1}$, and if $\Omega(X)\subseteq\mathcal{P}(X)$ meet-generates it, the other properties of Raney extensions are easy to check. If (2) holds, then $\mathcal{R}$-density by Lemma 5.9, and by the characterization in Proposition 3.2. By Proposition 3.8, and since every regular filter is exact, any Raney extension is $\mathcal{R}$-compact, thus (3) implies (4). If (4) holds, then (5) follows from the uniqueness result in Theorem 3.6. If (4) holds, then $\mathcal{U}(X)^{*}=\mathsf{Filt}_{\mathcal{R}}(\Omega(X))$, and so all neighborhood filters are regular. By Lemma 5.8, (1) follows. ∎
Let us then study the $T_{1}$ axiom more pointfreely. A space $X$ is $T_{1}$ if and only if $\mathcal{U}(X)$ is the same as the powerset $\mathcal{P}(X)$. Motivated by this, we define a Raney extension $(L,C)$ to be $T_{1}$ if and only if $C$ is a Boolean algebra.
Theorem 5.11.
For a frame $L$, the following are equivalent.
  1. 1.
    $L$ is subfit.
  2. 2.
    All exact filters of $L$ are regular.
  3. 3.
    $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ is a $T_{1}$ Raney extension.
  4. 4.
    There exists a $T_{1}$ Raney extension $(L,C)$.
  5. 5.
    There is a unique $T_{1}$ Raney extension on $L$, up to isomorphism. This is $(L,\mathsf{Filt}_{\mathcal{R}}(L)^{op})$.
Proof.
Suppose that $L$ is a subfit frame. By Proposition 2.14, all principal filters are regular filters. By Lemma 3.7, this implies that $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathsf{Filt}_{\mathcal{R}}(L)$. Now, suppose that $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathsf{Filt}_{\mathcal{R}}(L)$. This implies that $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})=(L,\mathsf{Filt}_{\mathcal{R}}(L)^{op})$, as regular filters are exact for every frame. By Proposition 2.12, the coframe $\mathsf{Filt}_{\mathcal{R}}(L)^{op}$ is a Boolean algebra. It is clear that condition (3) implies condition (4). Let us show that (4) implies (5). If $(L,B)$ is a Raney extension such that $B$ is Boolean, $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq B^{*}$ by Proposition 3.8. As $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathsf{Filt}(L)$ is dense, $B^{*}$ is dense, too, and as the only sublocale that is both Boolean and dense is the Booleanization this means $\mathsf{Filt}_{\mathcal{R}}(L)=\mathsf{Filt}_{\mathcal{E}}(L)=B^{*}$. Thus, $(L,B)$ and $(L,\mathsf{Filt}_{\mathcal{R}}(L)^{op})$ are isomorphic. Now, suppose that (5) holds. Then, $(L,\mathsf{Filt}_{\mathcal{R}}(L)^{op})$ is a Raney extension. This means that all principal filters are regular, and so by Proposition 2.14 the frame $L$ must be subfit. ∎
5.4 Scatteredness
The notion of scattered space is already present in classical topology, see for example [willard70]. In [simmons80] it is proven that a $T_{0}$ space is scattered if and only if $\mathsf{S}(\Omega(X))$ is Boolean. This motivates the definition of scattered frame (see [plewe00]): a frame is scattered if the coframe $\mathsf{S}(L)$ is Boolean. As proven in [ball16], a frame is scattered and subfit if and only if $\mathsf{S}(L)=\mathsf{S}_{\mathfrak{c}}(L)$. Subfit scattered frames are also fit, and so $\mathsf{S}(L)=\mathsf{S}_{\mathfrak{o}}(L)$.
Proposition 5.12.
For a subfit frame $L$, the following are equivalent.
  1. 1.
    The frame $L$ is scattered.
  2. 2.
    $\mathsf{Filt}_{\mathcal{SE}}(L)=\mathsf{Filt}_{\mathcal{E}}(L)=\mathsf{Filt}_{\mathcal{R}}(L)$.
  3. 3.
    $\mathsf{Filt}_{\mathcal{SE}}(L)=\mathsf{Filt}_{\mathcal{E}}(L)$.
  4. 4.
    The frame has a unique Raney extension, up to isomorphism.
  5. 5.
    $\mathsf{S}_{\mathfrak{o}}(L)=\mathsf{S}_{\mathfrak{c}}(L)$.
  6. 6.
    The frame has a unique Raney extension, up to isomorphism, and this is $(L,\mathsf{S}(L))$.
Proof.
Suppose that $L$ is a scattered subfit frame. Let $F$ be a strongly exact filter, by Theorem 2.8 this is $\{x\in L:S\subseteq\mathfrak{o}(x)\}$ for some sublocale $S$. By hypothesis, $S$ is a join $\bigvee_{i}\mathfrak{c}(x_{i})$ of closed sublocales, so that
$F=\{x\in L:x_{i}\vee x=1\text{ for all }i\in I\}=\bigcap_{i}\{x\in L:x\vee x_{i}=1\}.$
By the characterization of regular filters in Proposition 2.12, then, $\mathsf{Filt}_{\mathcal{SE}}(L)\subseteq\mathsf{Filt}_{\mathcal{R}}(L)$. This implies (2), as for all frames $\mathsf{Filt}_{\mathcal{R}}(L)\subseteq\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathsf{Filt}_{\mathcal{SE}}(L)$. It is clear that (2) implies (3). Let us show that (3) implies (4). The inclusion $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathsf{Filt}_{\mathcal{SE}}(L)$ holds for every frame. Now, suppose that in $L$ every strongly exact filter is exact. For any Raney extension $(L,C)$, we must have $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}\subseteq\mathsf{Filt}_{\mathcal{SE}}(L)$. Our assumption, then, implies $\mathsf{Filt}_{\mathcal{E}}(L)=C^{*}=\mathsf{Filt}_{\mathcal{SE}}(L)$. Suppose, now, that $L$ has a unique Raney extension, up to isomorphism. The pair $(L,\mathsf{S}_{\mathfrak{o}}(L))$ is a Raney extension. As $L$ is subfit, this must be a Boolean extension, by Theorem 5.11. As $\mathsf{S}_{\mathfrak{o}}(L)$ is a subcoframe of $\mathsf{S}(L)$, this means that in $\mathsf{S}(L)$ every fitted sublocale has a complement, which is itself a fitted sublocale. In particular, all joins of closed sublocales are fitted and so $\mathsf{S}_{\mathfrak{c}}(L)\subseteq\mathsf{S}_{\mathfrak{o}}(L)$. Finally, recall that the lattice $\mathsf{S}_{b}(L)$ of joins of complemented sublocales is $\mathsf{S}_{\mathfrak{c}}(L)$ for subfit frames. We then also have the reverse set inclusion $\mathsf{S}_{\mathfrak{o}}(L)\subseteq\mathsf{S}_{\mathfrak{c}}(L)$. If (5) holds, by subfitness $(L,\mathsf{S}_{\mathfrak{o}}(L))$ is a Boolean extension. Since this is the largest Raney extension, all its Raney extensions must be Boolean. By Theorem 5.11, when Boolean extensions exist, they are unique. Note also that $\mathsf{S}_{\mathfrak{o}}(L)=\mathsf{S}_{\mathfrak{c}}(L)$ implies that every closed sublocale is fitted, and this implies that the frame $L$ is subfit, hence $\mathsf{S}_{\mathfrak{o}}(L)=\mathsf{S}(L)$. Suppose, finally, that (6) holds. Because all subfit frames have a Boolean extension, by Theorem 5.11, $\mathsf{S}(L)$ must be Boolean, and so $L$ is scattered. ∎
6 Free and cofree constructions
6.1 Extensions of frame maps
We ask when a map $f:L\to M$ of frames can be extended to Raney extensions of these frames. In general, for Raney extensions $(L,C)$ and $(M,D)$, if such $\overline{f}:C\to D$ exists it has to be defined as $\overline{f}(c)=\bigwedge\{f(a):a\in{\uparrow}^{L}c\}$.
Theorem 6.1.
A frame map $f:L\to M$ extends to a map $\overline{f}:(L,C)\to(M,D)$ between Raney extensions if and only if $f^{-1}(F)\in C^{*}$ for every $F\in D^{*}$.
Proof.
First, suppose that there is a map $\overline{f}:(L,C)\to(M,D)$ of Raney extensions extending $f$. In particular, $\overline{f}$ preserves all meets and so it has a left adjoint $\overline{f}^{*}$. Consider $d\in D$. By adjointness, $f^{-1}({\uparrow}^{M}d)={\uparrow}^{L}\overline{f}^{*}(d)$. For the converse, suppose $f^{-1}(F)\in C^{*}$ for every $F\in D^{*}$. This means that for every $d\in D$ there is some $\overline{f}^{*}(d)\in C$ such that $f^{-1}({\uparrow}^{M}d)={\uparrow}^{L}\overline{f}^{*}(d)$, which is unique as $L$ meet-generates $C$. First, we claim that $\overline{f}^{*}:D\to C$ preserves all joins. To show $\bigvee_{i}\overline{f}^{*}(d_{i})=\overline{f}^{*}(\bigvee_{i}d_{i})$, it suffices to show ${\uparrow}^{L}\bigvee_{i}\overline{f}^{*}(d_{i})={\uparrow}^{L}\overline{f}^{*}(\bigvee_{i}d_{i})$.
${\uparrow}^{L}\bigvee_{i}\overline{f}^{*}(d_{i})$ $=\bigcap_{i}{\uparrow}^{L}\overline{f}^{*}(d_{i})=\bigcap_{i}f^{-1}({\uparrow}^{L}d_{i})=f^{-1}(\bigcap_{i}{\uparrow}^{L}d_{i})=f^{-1}({\uparrow}^{L}\bigvee_{i}d_{i})={\uparrow}^{L}\overline{f}^{*}(\bigvee_{i}d_{i}).$
Then, $\overline{f}^{*}:D\to C$ has a right adjoint, which we call $\overline{f}$. We claim that this is the required map. First, we show that it extends $f$. For $a\in L$,
$\overline{f}(a)$ =$=\bigvee\{d\in D:\overline{f}^{*}(d)\leq a\}=\bigvee\{d\in D:a\in{\uparrow}^{L}\overline{f}^{*}(d)\}$
=$=\bigvee\{d\in D:a\in f^{-1}({\uparrow}^{L}d)\}=\bigvee\{d\in D:d\leq f(a)\}=f(a).$
As it is a right adjoint, it preserves all meets. For finite joins, consider two arbitrary elements $\bigwedge_{i}a_{i},\bigwedge_{j}b_{j}\in C$ with $a_{i},b_{j}\in L$. Then
$\overline{f}(\bigwedge_{i}a_{i}\vee\bigwedge_{j}b_{j})=\bigwedge_{i,j}\overline{f}(a_{i}\vee b_{j})=\bigwedge_{i,j}\overline{f}(a_{i})\vee\overline{f}(b_{j})=\overline{f}(\bigwedge_{i}a_{i})\vee\overline{f}(\bigwedge_{j}b_{j}),$
where we have used that $\overline{f}$ preserves joins of $L$ as well as coframe distributivity. Then, the map $\overline{f}$ is the sought for map. ∎
In the coming subsections, we use this theorem to construct various free and cofree objects.
6.2 Free Raney extension on a frame
There is a forgetful functor $\pi_{1}:\mathbf{Raney}\rightarrow\mathbf{Frm}$ which forgets about the second component of the extension. We will show that this has a left adjoint.
Lemma 6.2.
Any frame morphism $f:L\rightarrow M$ extends to a Raney morphism
$f_{\mathcal{SE}}:(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})\rightarrow(M,\mathsf{Filt}_{\mathcal{SE}}(M)^{op}).$
Proof.
@@ 6.2 Free Raney extension on a frame
By Theorem 6.1, a frame morphism $f:L\rightarrow M$ lifts as required if preimages of strongly exact filters are strongly exact. Suppose, then, that $F\subseteq M$ is strongly exact. Suppose that the meet $\bigwedge_{i}x_{i}$ is strongly exact, and that $f(x_{i})\in F$. Because all frame morphisms preserve strongly exact meets, as well as strong exactness of meets, $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})\in F$, as desired. ∎
Then, there is a functor $\mathsf{Filt}_{\mathcal{SE}}:\mathbf{Frm}\to\mathbf{Raney}$ assigning to a frame $L$ the Raney extension $(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$, and to a morphism $f:L\to M$ the Raney map $f_{\mathcal{SE}}:(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})\to(M,\mathsf{Filt}_{\mathcal{SE}}(M)^{op})$ extending it, which exists by Lemma 6.2.
Theorem 6.3.
For a frame $L$, the pair $(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$ is the free Raney extension over it, that is, $\mathsf{Filt}_{\mathcal{SE}}\dashv\pi_{1}$. In particular, the category of frames is a full coreflective subcategory of $\mathbf{Raney}$.
Proof.
Suppose that $f:L\rightarrow M$ is a frame map. Let $(M,D)$ be a Raney extension. By Lemma 3.5, $D^{*}\subseteq\mathsf{Filt}_{\mathcal{SE}}(M)^{op}$, and by Lemma 6.2, preimages of strongly exact filters are strongly exact. Therefore, preimages of elements in $D^{*}$ are in $\mathsf{Filt}_{\mathcal{SE}}(L)$. By Theorem 6.1, the frame map lifts to a map $\overline{f}:(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})\to(M,D)$ extending $f$. ∎
6.3 Free spatial Raney extension on a spatial frame